Circular No. 5923

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994A IN UGC 8214 J. Mueller reports her discovery of a supernova of red mag ∼ 18 lo-

cated just to the north of a spiral arm in UGC 8214 (α = 13^{h}05^{m}.7, δ = +62^{o}29′, equinox 1950.0). SN 1994A, which was found on a plate taken by Mueller on Jan. 11 UT with the 1.2-m Oschin Telescope in the course of the second Palomar Sky Survey, is situated 3′′ east and 10′′.1 south of the galaxy nucleus. N. Reid notes that a red Survey plate taken on 1987 May 23 by J. Phinney and himself shows no object at the position of SN 1994A. A. V. Filippenko and T. Matheson, University of California at Berkeley, report that a CCD spectrogram (range 310-1000 nm) obtained on Jan. 19 with the 3-m Shane reflector at Lick Observatory shows that the supernova is of type II; broad Hα emission is visible, but the corresponding absorption component is weak or absent.

SUPERNOVA 1994B IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

supernova (V = 19.0) located at α = 8^{h}17^{m}51^{s}.64, δ = +15^{o}53′20′′.5 (equinox 1950.0), which is 4′′.2 east and 3′′.9 north of the center of the face-on spiral host galaxy (V = 17.5). SN 1994B was found on films taken on Jan. 16.98 and 17.94 UT with the OCA 0.9-m Schmidt telescope. A nearby star (V = 17.5) has end figures 52^{s}.85, 54′28′′.7. Filippenko and Matheson report that an uncalibrated CCD spectrogram, obtained as above on Jan. 19 UT, suggests that SN 1994B is a type-Ia supernova, near maximum brightness. A broad Si II absorption trough is visible. The object is superposed on an H II region, from which an approximate redshift of 0.089 was derived for the parent galaxy. Photometry is encouraged.

GRO J0422+32 P. Zhao, P. Callanan, M. Garcia, and J. McClintock, Smithsonian As-

trophysical Observatory, report additional photometry (cf. IAUC 5907): 1993 Dec. 16.09 UT, R = 14.89, V = 15.18; 23.30, 15.45, –; 1994 Jan. 4.15, 18.03, 18.76; 4.40, 18.19, –; 5.09, 18.76, –; 5.41, 19.13, –; 6.17, 18.88, 19.52; 6.35, 18.59, 19.00; 7.26, 17.77, 18.10; 8.31, 17.31, 17.68; 9.20, 17.00, 17.39; 10.29, 17.12, –; 11.29, 17.25, –; 12.29, 17.53, –; 13.29, 18.02, –; 14.16, 18.40, 19.11; 16.17, 17.97, 18.32; 17.13, 17.89, 18.40. During the past month, maxima have occurred on Dec. 10 and Jan. 9 and 16, with minima on Jan. 5 and 14; each cycle is shorter and has a smaller amplitude.

1994 January 19 Daniel W. E. Green